A concerned citizen group claims that 50% of the people in the city A support making beers illegal. Yoi decide to test this claim and ask a random sample of 315 out of 6000 people in the concerned city whether they support making beers illegal, 49% support this law. At 0.05 significance level, is there enough evidence to support this claim?

1
Expert's answer
2022-04-27T12:58:14-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.5"

"H_1:p\not=0.5"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}"

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.49-0.5}{\sqrt{\dfrac{0.5(1-0.5)}{315}}}\approx-0.3550"

Since it is observed that "|z| = 0.355 \le 1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p =2P(Z<-0.3550)= 0.72259," and since "p = 0.7226 \ge 0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than 0.5, at the "\alpha = 0.05" significance level.